OpenStudy (anonymous):

Rate of Change , you are given the dollar value of a product in 2004 and the rate at which the value of the product is expected to change during the next 5 years. Write a linear equation that gives the dollar value (v) of the product in terms of the year (Let t=0 represent 2000.) 2004 value= $156 with $4.50 increase per year

OpenStudy (dumbcow):

when t=4, v = 156 the equation of linear line y = mx+b m=slope b = y-intercept or initial value apply it to this problem v = 4.5t + b to solve for b, use given info (v=156 when t=4) 156 = 4.5(4) + b b = 156 - 4.5(4) b = 156 - 18 b = 138 Linear equation: v = 4.5t +138 this means the product had value of $138 in 2000

OpenStudy (anonymous):

thanks!!! :D

OpenStudy (dumbcow):

your welcome